OpenStudy (anonymous):

39 - 22 1/6.. can u please show me the steps

OpenStudy (anonymous):

is the answer 22 1/6

OpenStudy (anonymous):

sorry 17 1/6

OpenStudy (anonymous):

should be 16 5/6

OpenStudy (anonymous):

can u please show me how to do it

OpenStudy (anonymous):

since the 39 is a whole number, you can change it to 38 6/6 for this equation. Subtracting 38 and 22 will give you the 16, and subtracting 1/6 from 6/6 will give you 5/6. So you're left with 16 5/6

OpenStudy (anonymous):

how do i change 39 to 38 6/6

OpenStudy (anonymous):

it is the same number. 6/6 = 1

OpenStudy (anonymous):

\[39 - 22\frac{1}{6} = 39 - \frac{133}{6} = \frac{234-133}{6} = \frac{101}{6} = 16\frac{5}{6}\]

OpenStudy (anonymous):

where does the 234 comes from??? sorry but lost

OpenStudy (anonymous):

We did it two different ways.

OpenStudy (anonymous):

i see what u did but i dont know where 234 came from

OpenStudy (anonymous):

the 234 comes from changing the whole number (39) to a fraction since 39 as a whole is the same as 39/1, to make it equal to the other fractions, you need to make the denominator (bottom number) the same. so you would multiply 39/1 by 6/6

OpenStudy (anonymous):

133 is figured out the same way. You have -22 1/6. to make it all a fraction, you would multiply the -22 by 6/6 to get -132/6, then add it to the 1/6 you already have to get -133/6

OpenStudy (anonymous):

thanks a lot..

OpenStudy (anonymous):

Does it make sense?

OpenStudy (anonymous):

yes it does

OpenStudy (anonymous):

sorry.. 6 - 2 3/8.. is the answer to that 4 5/8

OpenStudy (anonymous):

That is wrong

OpenStudy (anonymous):

Remember that the 3/8 needs to have a fraction to subract from

OpenStudy (anonymous):

so to do that, you need to make 6 have a fraction by changing one of the whole numbers into a fraction to get 5 8/8, which is still the same as 6

OpenStudy (anonymous):

what i did was making the 6: (5 8/8/) and then i substracted the 2 3/8 giving me a 3 5/8

OpenStudy (anonymous):

you did it right that time. I think you typed it wrong the first time

OpenStudy (anonymous):

ok.. thankssss .. i think i got it now

OpenStudy (anonymous):

No problem, glad to help